OpenStudy (anonymous):

Plzzzz help will give medal and fan The lengths of turtles in a park are normally distributed with a mean of 25.9 inches and a standard deviation of 6.3 inches. What is the percentile rank of a turtle whose length is 21 inches? 22% 19% 30% 10%

OpenStudy (anonymous):

ok so first what you would need to do is kinda like make a picture to help you

OpenStudy (anonymous):

then what you would

OpenStudy (anonymous):

ok

OpenStudy (tylerd):

sec trying to solve first

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

z = (21 - 25.9)/6.3 = -0.7778 P(x < = 21) = P(z < = - 0.7778) = 0.22 = 22% (Ans.)

OpenStudy (anonymous):

are u sure

OpenStudy (anonymous):

100% positive

OpenStudy (anonymous):

Tyler wat do u think?

OpenStudy (tylerd):

shes probably right but i wouldnt be able to explain it

OpenStudy (anonymous):

ok thanks for help guys

OpenStudy (anonymous):

welcome! @NightMath

OpenStudy (anonymous):

thanks for the medal and fan!

OpenStudy (anonymous):

mp